Given the following raw data, make a measures of central tendency table ( Class Interval, Tally, Frequency, Class Mark, Class Boundaries, <cmf, >cmf, rf, %rf ). Determine Mean, Median, and Mode, then find with Q1, Q2, and Q3 with complete solution. Test scores obtained by sixty students in a statistics class. 48 73 57 57 69 88 11 80 82 47 46 70 49 45 75 81 33 65 38 59 94 59 62 36 58 69 45 55 58 65 30 49 73 29 41 53 37 35 61 48 22 51 56 55 60 37 56 59 57 36 12 36 50 63 68 30 56 70 53 28
1
Expert's answer
2021-10-21T07:32:37-0400

Least to Greatest order:

11, 12, 22, 28, 29, 30, 30, 33, 35, 36, 36, 36, 37, 37, 38, 41, 45, 45, 46, 47, 48, 48, 49, 49, 50, 51, 53, 53, 55, 55, 56, 56, 56, 57, 57, 57, 58, 58, 59, 59, 59, 60, 61, 62, 63, 65, 65, 68, 69, 69, 70, 70, 73, 73, 75, 80, 81, 82, 88, 94


Class interval:10

"n=60"


"mean=Q_2=\sum x_i/n=53.08"

"median=(x_{30}+x_{31})\/2=(55+56)\/2=55.5"

Mode: 36,56,57,59 (appears 3 times)

"Q_1=(x_{15}+x_{16})\/2=(38+41)\/2=39.5"

"Q_3=(x_{45}+x_{46})\/2=(63+65)\/2=64"